OpenStudy (anonymous):

A student takes out two loans totaling $10,000 to help pay for college expenses. One loan is at 4% simple interest, and the other is at 9% simple interest. The first-year interest is . Find the amount of the loan at 9%. A) $3000 B) $120 C) $630 D) $7000

OpenStudy (anonymous):

I think we are missing the value of the first year interest. Do you have that number?

OpenStudy (anonymous):

$750

OpenStudy (anonymous):

We know that there are two loans. We know that the sum of the two loans is $10,000. So let's assign a variable to each loan. Let a= loan one Let b=loan two Since we have two variables, we need two equations to solve this problem. The first equation is: loan one + loan two = $10,000 or a+b=10,000

OpenStudy (anonymous):

We also know that the interest charge on loan one plus the interest charge on loan two is $750. So the second equation looks like this: interest charge loan one + interest charge loan two = 750 To find out the interest charge in simple interest on each loan, we change the percent to a decimal and then multiply the decimal by the loan amount. So for loan one, change 4% to .04. Next multiply by the loan amount, which we are using the variable a to represent. So the interest charge for loan one looks like this: .04a Do the same thing for loan two. Change 9% to .09 decimal. Next multiply by the loan amount, which we are using variable b to represent. So the interest charge for loan two looks like this: .09b We can now write our second equation with the variables. Our second equation looks like this: .04a+.09b=750

OpenStudy (anonymous):

We now have two equations. So we can solve for the two variables. .04a+.09b=750 a+b=10,000 Rearrange the bottom equation. a=10,000-b Substitute into the top equation. .04(10,000-b)+.09b=750 Distribute .04(10,000-b) 400-.04b+.09b=750 Combine like terms. 400+05b=750 rearrange the equation. .05b=750-400 Solve for b .05b=350 b=7,000 So the amount of the second loan is 7,000 dollars.

OpenStudy (anonymous):

Thank you so much!